OpenStudy (anonymous):

(5,-10),m=1/5 written in slope intercept form

OpenStudy (anonymous):

y=1/5x - 11

OpenStudy (anonymous):

slope intercept form: y - y1 = m(x - x1). Fill in the numbers, y - (-10) = 1/5(x - 5) --> y+10=0.2x - 1 --> y = 0.2 - 11.